# Waring and Waring–Goldbach subbases with prescribed representation function

### v1

@@ Abstract.
We study representations of integers as sums of $h$ elements from subsets of $k$-th powers $\mathbb{N}^{k}$ and $k$-th powers of primes $\mathbb{P}^{k}$ for $k\geq 1$. Extending work by Vu, Wooley, and others, we show that for $h\geq h_{k}=O(8^{k}k^{2})$, if $F$ is a regularly varying function satisfying $\lim_{n\to\infty}F(n)/\log n=\infty$ and certain growth conditions, then there exists $A\subseteq\mathbb{N}^{k}$ such that $r_{A,h}(n)\sim\mathfrak{S}_{k,h}(n)F(n)$, where $\mathfrak{S}_{k,h}(n)$ is the singular series from Waring’s problem.
For $h\geq 2^{k}+1$, we show the existence of $A\subseteq\mathbb{P}^{k}$ with $r_{A,h}(n)\sim\mathfrak{S}^{*}_{k,h}(n)cn^{\kappa}$ for $10<\kappa<h/k-$ 1 and $c>0$, where $\mathfrak{S}^{*}_{k,h}(n)$ is the singular series from Waring–Goldbach’s problem. Additionally, for $10\leq\kappa\leq h/k-$ 1 and functions $\psi$ satisfying $\psi(x)\asymp_{\lambda}\psi(x^{\lambda})$ for every $\lambda>0$, if $\log x\ll x^{\kappa}\psi(x)\ll x^{h/k-1}/(\log x)^{h}$, then there exists $A\subseteq\mathbb{P}^{k}$ with $r_{A,h}(n)\asymp n^{\kappa}\psi(n)$ for $n$ satisfying certain congruence conditions, producing thin subbases of prime powers when $\kappa=0$, $\psi=\log$.
@@ 1.1. Waring subbases
    $\lim_{x\to\infty}\frac{F(x)}{\log x}=\infty,\qquad F(x)\leq(1+o(1))\frac{\Gamma(1+1/k)^{h}}{\Gamma(h/k)}x^{h/k-1},$
    such that $F(x)/x^{h/k-1}$ is asymptotic to a non-increasing function. Then, there exists a subset $A\subseteq\mathbb{N}^{k}$ such that
@@
Remark.
In Theorem 1.1 (ii), if $F(x)\ll x^{h/k-1-\varepsilon}$ for some $\varepsilon>0$, then $F(x)/x^{h/k-1}$ is necessarily asymptotic to a non-increasing function, since $F(x)/x^{h/k-1}\sim\sup_{y\geq x}F(y)/y^{h/k-1}$ (cf. BGT [1, Theorem 1.5.3]).
@@ 1.2. Waring–Goldbach subbases
Let $k\geq 1$ be an integer, and $h\geq 2^{k}+1$. For every $\omega\geq 1/h$, we have
@@
Let $k\geq 1$ be an integer, $h\geq 2^{k}+1$, and suppose that $n\equiv h\leavevmode\nobreak\ (\mathrm{mod}\leavevmode\nobreak\ K(k))$.
@@ 1.3. General setup
    (Counting solutions with weights) There exists $H_{\textnormal{(iii)}}\in\mathbb{Z}_{\geq 1}$ such that the following holds: Let $h\geq H_{\textnormal{(iii)}}$, and suppose $F(n)\leq(1+o(1))r_{B,h}(n)$ is such that $f(x):=(xF(x))^{1/h}=x^{\omega}\phi(x)$ satisfies $1/h\leq\omega\leq\beta$, with $f(x)/B(x)$ asymptotic to a non-increasing function. Then:
@@
Condition (ii) asks for $hB$ to have almost as low additive energy as possible, in the sense that $\sum_{n\leq x}r_{B,h}(n)^{2}\geq(\sum_{n\leq x}r_{B,h}(n))^{2}/x\asymp B(x)^{2h}/x$ by Cauchy–Schwarz. It will only be used in the proof of Proposition 2.2. In condition (iii), we roughly ask for the expectation of $r_{\mathscr{A},h}(n)$ for a random subset $\mathscr{A}\subseteq B$ of density $|\mathscr{A}\cap[1,x]|\approx f(x)$ to have a certain asymptotic behaviour, over a subset $\mathscr{S}\subseteq\mathbb{Z}_{\geq 0}$ that is usually an arithmetic progression. We observe that if $\omega<\beta$, then $f(x)/B(x)$ is necessarily asymptotic to a non-increasing function since $f(x)/B(x)\sim\sup_{y\geq x}f(y)/B(y)$ by [1, Theorem 1.5.3]. The only place we need to assume $f(x)/B(x)$ monotone is Proposition 2.2.
@@
Moreover, if $F=\log$, then there exists $A\subseteq B$ such that
$r_{A,h}(n)\asymp\log n\quad\text{for}\quad n\in\mathscr{S}.$
@@ 2. Random subsets
be a regularly varying function such that $cf(x)\leq B(x)$ and $1/h\leq\omega\leq\beta$. We suppose that $f(x)/B(x)$ is non-increasing. [^2] Since $f(x)/B(x)$ is asymptotic to a non-increasing function by condition (iii), and we only care about asymptotics, we may substitute $f(x)$ by $B(x)(\sup_{y>x}f(y)/B(y))$ if needed. Now consider a probability space whose elements are subsets $0\in\mathscr{A}\subseteq B$, where
@@ 2.1. Expectation of $r_{\mathscr{A},h}(n)$
For $\alpha\in[0,1)$, $x\in\mathbb{R}_{\geq 0}$, let
@@
The function $f(x)/B(x)$ is non-increasing, so we have $\sum_{x/2^{j+1}<n\leq x/2^{j}}\big{|}\frac{f(n)}{B(n)}-\frac{f(n+1)}{B(n+1)}\big{|}\ll f(2^{-j}x)/B(2^{-j}x)$. It follows that
@@
It is sufficient to work with $f(x)$ under the conditions
$\frac{f(x)}{(x\log x)^{1/h}}\xrightarrow{x\to\infty}\infty,\quad\text{and}\quad cf(x)\leq x.$
@@ 3.4. Disjoint families of representations
where we set $\lambda:=32hK\log n$. If $F(n)/\log n\to\infty$ as $x\to\infty$, then Lemma 3.3 implies $(\lambda\mathbb{E}(\rho^{(\delta\text{-normal})}_{\mathscr{A},h}(n)))^{1/2}=o(\mathbb{E}(\rho^{(\delta\text{-normal})}_{\mathscr{A},h}(n)))$. Thus, by the Borel–Cantelli lemma, we have $\rho^{(\delta\text{-normal})}_{\mathscr{A},h}(n)\stackrel{{\scriptstyle\text{a.s.}}}{{\sim}}\mathbb{E}(\rho^{(\delta\text{-normal})}_{\mathscr{A},h}(n))$. For $F=c\log$ with $c>1$ chosen sufficiently large to ensure $\mathbb{E}(\rho_{\mathscr{A},h}(n))>32hK\log n$, it follows that $\rho^{(\delta\text{-normal})}_{\mathscr{A},h}(n)\stackrel{{\scriptstyle\text{a.s.}}}{{\asymp}}\log n$.
@@ 4.1. Waring satisfies condition (iii)
Fix $k\geq 1$ an integer, $h\geq 2^{k}+1$, let $\omega\geq 1/h$ be a real number, and $N$ be a large integer. For the next two sections, we will deal with $\mathbb{P}^{k}$. Define
@@ 5.1.1. Lemmas
$=\frac{P}{\varphi(q)}\frac{x^{\omega}}{\omega}+\int_{2^{-}}^{x}\frac{t^{\omega}}{t^{1/k}/\log t}\,\mathrm{d}\bigg{(}\pi(t^{1/k};q,r_{i})-\frac{\mathrm{li}(t^{1/k})}{\varphi(q)}\bigg{)}.$
@@
$\mathfrak{S}_{k,h}^{*}(N,Q)=\bigg{(}1+O_{\varepsilon}\bigg{(}\frac{1}{Q^{1/2^{k-1}-\varepsilon}}\bigg{)}\bigg{)}\mathfrak{S}_{k,h}^{*}(N)\quad\text{as}\quad Q\to\infty.$
@@
By Lemmas 8.10 and 8.12 in Hua [8], we have $\mathfrak{S}_{k,\ell}^{*}(N)\gg 1$ for $N\equiv h\leavevmode\nobreak\ (\mathrm{mod}\leavevmode\nobreak\ K(k))$ whenever $\ell\geq 3$ for $k=1$, $\ell\geq 5$ for $k=2$, and $\ell\geq 3k$ for $k\geq 3$. In all cases, this holds true for $h\geq 2^{k}+1$.
For the second part, let $q\geq 1$, $1\leq a\leq q$ with $(a,q)=1$. By Weyl’s inequality we have $S(q,a)\ll_{\varepsilon}q^{1-1/2^{k-1}+\varepsilon}$ uniformly in $q$ (cf. Nathanson [12, Theorem 4.4]). Thus, since $\varphi(q)\gg_{\varepsilon}q^{1-\varepsilon}$ and $h\geq 2^{k}+1$, we have
@@
$\ll_{\varepsilon}\sum_{q>Q}(q^{-1/2^{k-1}+\varepsilon})^{h}\,q$
$\ll_{\varepsilon}\sum_{q>Q}q^{-1-1/2^{k-1}+\varepsilon}\asymp Q^{-1/2^{k-1}+\varepsilon}.\qed$
@@ 5.1.2. Integral over major arcs
By (5.4), and the fact that $\sum_{n\leq t}n^{\omega-1}=\frac{t^{\omega}}{\omega}+O(1)$, we have $\sum_{n\leq t}D(n)=O_{A}(t/(\log t)^{A})$. By using partial summation we conclude, since $|\vartheta|\leq Q/N=(\log N)^{C}/N$, that
@@
$=\bigg{(}1+O_{\varepsilon}\bigg{(}\frac{1}{(\log N)^{C/2^{k-1}-\varepsilon}}\bigg{)}\bigg{)}\mathfrak{S}^{*}_{k,h}(N)\int_{0}^{1}\widetilde{u}(\vartheta;N)^{j}\,u(\vartheta;N)^{h-j}\,e(-N\vartheta)\,\mathrm{d}\vartheta$
@@
$=\bigg{(}1+O_{\varepsilon}\bigg{(}\frac{1}{(\log N)^{C/2^{k-1}-\varepsilon}}\bigg{)}\bigg{)}\mathfrak{S}^{*}_{k,h}(N)\int_{-Q/N}^{Q/N}\widetilde{u}(\vartheta;N)^{j}u(\vartheta;N)^{h-j}\,e(-N\vartheta)\,\mathrm{d}\vartheta$
@@ 5.2. Minor arcs
$\bigg{|}\sum_{j=1}^{h}\int_{\mathfrak{m}}\widetilde{T}(\alpha;N)^{j}\,T(\alpha;N)^{h-j}\,e(-N\alpha)\,\mathrm{d}\alpha\bigg{|}\ll_{R}\frac{N^{h\omega-1}}{(\log N)^{R}},$
@@ 5.2.1. Lemmas
The first lemma we need is equivalent to showing that $\mathbb{P}^{k}$ satisfies condition (ii) of Theorem 1.4 with $H_{\textnormal{(ii)}}\leq 2^{k-1}$.
@@
There exists $c(k)\in\mathbb{R}$ such that, for every $\ell\geq 2^{k-1}$,
$\sum_{n\leq x}r_{\mathbb{P}^{k},\ell}(n)^{2}\leq x^{2\ell/k-1}(\log x)^{c(k)}.$
@@
$\sum_{n\leq x}r_{\mathbb{P}^{k},2^{k-1}}(n)^{2}\leq\int_{0}^{1}|g(\alpha;x)|^{2^{k}}\,\mathrm{d}\alpha\leq\sum_{n\leq 2^{k-1}x}r_{\mathbb{P}^{k},2^{k-1}}(n)^{2},$
@@
$n_{1}+\cdots+n_{2^{k-1}}=m_{1}+\cdots+m_{2^{k-1}}$ (5.8)
with $n_{i},m_{j}\leq x$ and $n_{i},m_{j}\in\mathbb{P}^{k}$. By Hua’s inequality (cf. Hua [8, Theorem 4]), there exists $c(k)\in\mathbb{R}$ such that
@@
$\ll x^{2\ell/k-1}(\log x)^{c(k)}.\qed$

### v2

@@ Abstract.
We study $r_{A,h}(n)$, the number of representations of integers $n$ as sums of $h$ elements from subsets $A$ of $k$-th powers $\mathbb{N}^{k}$ and $k$-th powers of primes $\mathbb{P}^{k}$ for $k\geq 1$. Extending work by Vu, Wooley, and others, we show that for $h\geq h_{k}=O(8^{k}k^{2})$, if $F$ is a regularly varying function satisfying $\lim_{n\to\infty}F(n)/\log n=\infty$, then there exists $A\subseteq\mathbb{N}^{k}$ such that
$r_{A,h}(n)\sim\mathfrak{S}_{k,h}(n)F(n),$
where $\mathfrak{S}_{k,h}(n)$ is the singular series from Waring’s problem.
For $h\geq 2k^{2}(2\log k+\log\log k+O(1))$, we show the existence of $A\subseteq\mathbb{P}^{k}$ with
$r_{A,h}(n)\sim\mathfrak{S}^{*}_{k,h}(n)cn^{\kappa}$
for $10<\kappa<h/k-$ 1 and $c>0$, where $\mathfrak{S}^{*}_{k,h}(n)$ is the singular series from Waring–Goldbach’s problem. Additionally, for $10\leq\kappa\leq h/k-$ 1 and functions $\psi$ satisfying $\psi(x)\asymp_{\lambda}\psi(x^{\lambda})$ for every $\lambda>0$, if $\log x\ll x^{\kappa}\psi(x)\ll x^{h/k-1}/(\log x)^{h}$, then there exists $A\subseteq\mathbb{P}^{k}$ with $r_{A,h}(n)\asymp n^{\kappa}\psi(n)$ for $n$ satisfying certain congruence conditions, producing thin subbases of prime powers when $\kappa=0$, $\psi=\log$.
@@ 1.1. Waring subbases
    $\lim_{x\to\infty}\frac{F(x)}{\log x}=\infty,\qquad F(x)\leq(1+o(1))\frac{\Gamma(1+1/k)^{h}}{\Gamma(h/k)}x^{h/k-1}.$
    Then, there exists a subset $A\subseteq\mathbb{N}^{k}$ such that
@@ 1.2. Waring–Goldbach subbases
Let $k\geq 1$ be an integer, and
$h\geq h_{k}^{*}:=\begin{cases}2^{k}+1&\text{ if }1\leq k\leq 11,\\ \lceil 2k^{2}(2\log k+\log\log k+2.5)\rceil&\text{ if }k\geq 12.\end{cases}$ (1.4)
For every $\omega\geq 1/h$, we have
@@
Let $k\geq 1$ be an integer, $h\geq h_{k}^{*}$ where $h_{k}^{*}$ is as in (1.4), and suppose that $n\equiv h\leavevmode\nobreak\ (\mathrm{mod}\leavevmode\nobreak\ K(k))$.
@@ 1.3. General setup
    (Counting solutions with weights) There exists $H_{\textnormal{(iii)}}\in\mathbb{Z}_{\geq 1}$ such that the following holds: Let $h\geq H_{\textnormal{(iii)}}$, and suppose $F(n)\leq(1+o(1))r_{B,h}(n)$ is such that $f(x):=(xF(x))^{1/h}=x^{\omega}\phi(x)$ satisfies $1/h\leq\omega\leq\beta$. Then:
@@
Condition (ii) asks for $hB$ to have almost as low additive energy as possible, in the sense that $\sum_{n\leq x}r_{B,h}(n)^{2}\geq(\sum_{n\leq x}r_{B,h}(n))^{2}/x\asymp B(x)^{2h}/x$ by Cauchy–Schwarz. It will only be used in the proof of Proposition 2.2. In condition (iii), we roughly ask for the expectation of $r_{\mathscr{A},h}(n)$ for a random subset $\mathscr{A}\subseteq B$ of density $|\mathscr{A}\cap[1,x]|\approx f(x)$ to have a certain asymptotic behaviour, over a subset $\mathscr{S}\subseteq\mathbb{Z}_{\geq 0}$ that is usually an arithmetic progression.
@@
If $F(n)\gg\log n$, then there exists $A\subseteq B$ such that
$r_{A,h}(n)\asymp F(n)\quad\text{for}\quad n\in\mathscr{S}.$
@@ 2. Random subsets
be a regularly varying function such that $cf(x)\leq B(x)$ and $1/h\leq\omega\leq\beta$. Now consider a probability space whose elements are subsets $0\in\mathscr{A}\subseteq B$, where
@@ 2.1. Expectation of $r_{\mathscr{A},h}(n)$
Since
$\sum_{\begin{subarray}{c}x_{1},\ldots,x_{\ell}\in B\\ c_{1}x_{1}+\cdots+c_{\ell}x_{\ell}=n\end{subarray}}\frac{f(x_{1})}{B(x_{1})}\cdots\frac{f(x_{\ell})}{B(x_{\ell})}\asymp\sum_{\begin{subarray}{c}x_{1},\ldots,x_{\ell}\in B\\ c_{1}x_{1}+\cdots+c_{\ell}x_{\ell}=n\end{subarray}}G(x_{1})\cdots G(x_{\ell})$
for every $G(x)\asymp f(x)/B(x)$, we may substitute $f$ by some function asymptotic to $f$ before proceeding with the calculations. Write $f(x)=x^{\omega}\phi(x)$ and $f(x)/B(x)=x^{\omega-\beta}\vartheta(x)$.
  • If $\omega<\beta$, substitute $f$ by $f(x)\sim B(x)\sup_{y\geq x}f(y)/B(y)$ (cf. BGT [1, Theorem 1.5.3]), so that $f(x)/B(x)$ is non-increasing.
  • If $\omega=\beta$, substitute $f$ by $f(x)\sim f(x)\vartheta(x)/\xi(x)$, where
    $\xi(x):=\sup_{\begin{subarray}{c}2^{J}\leq y<2^{J+1}\\ J=\lfloor\frac{\log x}{\log 2}\rfloor\end{subarray}}\vartheta(y).$
    Since $\vartheta$ is of slow variation, we have $\lim_{x\to\infty}\sup_{\lambda\in[\frac{1}{2},2]}\vartheta(\lambda x)/\vartheta(x)=1$ by uniform convergence, therefore $\vartheta(x)\sim\xi(x)$.
For $f$ like described above, we have the following:
Lemma 2.5.
$\sum_{x\leq n\leq 2x}\bigg{|}\frac{f(n)}{B(n)}-\frac{f(n+1)}{B(n+1)}\bigg{|}\ll\frac{f(x)}{B(x)}$
Proof.
If $\omega<\beta$, then $f(x)/B(x)$ is non-increasing, so the result follows since $f(x)/B(x)\asymp f(2x)/B(2x)$. If $\omega=\beta$, then we have a sum of the form
$\sum_{x\leq n\leq 2x}|\xi(n)-\xi(n+1)|$
where $\xi$ is constant in the intervals $[2^{J},2^{J+1})$, $J\in\mathbb{Z}_{\geq 0}$. This implies that for $n\in[x,2x]$, $\xi(n)=\xi(n+1)$ for all except possibly one $n$. Therefore
$\sum_{x\leq n\leq 2x}|\xi(n)-\xi(n+1)|\leq\sup_{x\leq y\leq 2x}|\xi(y)-\xi(y+1)|\ll\xi(x)=\frac{f(x)}{B(x)}.\qed$
Now, for $\alpha\in[0,1)$, $x\in\mathbb{R}_{\geq 0}$, let
@@
By Lemma 2.5, we have $\sum_{x/2^{j+1}<n\leq x/2^{j}}\big{|}\frac{f(n)}{B(n)}-\frac{f(n+1)}{B(n+1)}\big{|}\ll f(2^{-j}x)/B(2^{-j}x)$. It follows that
@@
It is sufficient to work with $f(x)$ such that
$f(x)\gg(x\log x)^{1/h},\quad\text{and}\quad cf(x)\leq x.$
@@ 3.4. Disjoint families of representations
where we set $\lambda:=32hK\log n$. If $F(n)/\log n\to\infty$ as $x\to\infty$, then Lemma 3.3 implies $(\lambda\mathbb{E}(\rho^{(\delta\text{-normal})}_{\mathscr{A},h}(n)))^{1/2}=o(\mathbb{E}(\rho^{(\delta\text{-normal})}_{\mathscr{A},h}(n)))$. Thus, by the Borel–Cantelli lemma, we have $\rho^{(\delta\text{-normal})}_{\mathscr{A},h}(n)\stackrel{{\scriptstyle\text{a.s.}}}{{\sim}}\mathbb{E}(\rho^{(\delta\text{-normal})}_{\mathscr{A},h}(n))$. For $F\gg\log$, we can choose $c>1$ sufficiently large to ensure that $\mathbb{E}(\rho_{\mathscr{A},h}(n))\gg c^{h}F(n)>32hK\log n$, so that it follows that $\rho^{(\delta\text{-normal})}_{\mathscr{A},h}(n)\stackrel{{\scriptstyle\text{a.s.}}}{{\asymp}}F(n)$.
@@ 4.1. Waring satisfies condition (iii)
For the next two sections, we will deal with $\mathbb{P}^{k}$. Fix $k\geq 1$ an integer, and let
$h_{k}^{*}:=\begin{cases}2^{k}+1&\text{ if }1\leq k\leq 11,\\ \lceil 2k^{2}(2\log k+\log\log k+2.5)\rceil&\text{ if }k\geq 12.\end{cases}$
Fix $h\geq h^{*}_{k}$ an integer, let $\omega\geq 1/h$ be a real number, and $N$ be a large integer. Define
@@ 5.1.1. Lemmas
$=\frac{P}{\varphi(q)}\frac{x^{\omega}}{\omega}+\sum_{i=1}^{P}\int_{2^{-}}^{x}\frac{t^{\omega}}{t^{1/k}/\log t}\,\mathrm{d}\bigg{(}\pi(t^{1/k};q,r_{i})-\frac{\mathrm{li}(t^{1/k})}{\varphi(q)}\bigg{)}.$
@@
$\mathfrak{S}_{k,h}^{*}(N,Q)=\bigg{(}1+O_{\varepsilon}\bigg{(}\frac{1}{Q^{1/k-\varepsilon}}\bigg{)}\bigg{)}\mathfrak{S}_{k,h}^{*}(N)\quad\text{as}\quad Q\to\infty.$
@@
By Lemmas 8.10 and 8.12 in Hua [8], we have $\mathfrak{S}_{k,\ell}^{*}(N)\gg 1$ for $N\equiv h\leavevmode\nobreak\ (\mathrm{mod}\leavevmode\nobreak\ K(k))$ whenever $\ell\geq 3$ for $k=1$, $\ell\geq 5$ for $k=2$, and $\ell\geq 3k$ for $k\geq 3$. In all cases, this holds true for $h\geq h_{k}^{*}$.
For the second part, let $q\geq 1$, $1\leq a\leq q$ with $(a,q)=1$. By [8, Theorem 1], we have $S(q,a)\ll_{\varepsilon}q^{1-1/k+\varepsilon}$ uniformly in $q$. Thus, since $\varphi(q)\gg_{\varepsilon}q^{1-\varepsilon}$ and $h\geq h^{*}_{k}\geq 2k+1$, we have
@@
$\ll_{\varepsilon}\sum_{q>Q}(q^{-1/k+\varepsilon})^{h}\,q$
$\ll_{\varepsilon}\sum_{q>Q}q^{-1-1/k+\varepsilon}\asymp Q^{-1/k+\varepsilon}.\qed$
@@ 5.1.2. Integral over major arcs
By (5.4), and the fact that $\sum_{n\leq t}n^{\omega-1}=\frac{t^{\omega}}{\omega}+O(1)$, we have $\sum_{n\leq t}D(n)=O_{R}(t/(\log t)^{R})$. By using partial summation we conclude, since $|\vartheta|\leq Q/N=(\log N)^{C}/N$, that
@@
$=\bigg{(}1+O_{\varepsilon}\bigg{(}\frac{1}{(\log N)^{C/k-\varepsilon}}\bigg{)}\bigg{)}\mathfrak{S}^{*}_{k,h}(N)\int_{0}^{1}\widetilde{u}(\vartheta;N)^{j}\,u(\vartheta;N)^{h-j}\,e(-N\vartheta)\,\mathrm{d}\vartheta$
@@
$=\bigg{(}1+O_{\varepsilon}\bigg{(}\frac{1}{(\log N)^{C/k-\varepsilon}}\bigg{)}\bigg{)}\mathfrak{S}^{*}_{k,h}(N)\int_{-Q/N}^{Q/N}\widetilde{u}(\vartheta;N)^{j}u(\vartheta;N)^{h-j}\,e(-N\vartheta)\,\mathrm{d}\vartheta$
@@ 5.2. Minor arcs
$\sum_{j=1}^{h}\bigg{|}\int_{\mathfrak{m}}\widetilde{T}(\alpha;N)^{j}\,T(\alpha;N)^{h-j}\,e(-N\alpha)\,\mathrm{d}\alpha\bigg{|}\ll_{R}\frac{N^{h\omega-1}}{(\log N)^{R}},$
@@ 5.2.1. Lemmas
The first lemma we need is equivalent to showing that $\mathbb{P}^{k}$ satisfies condition (ii) of Theorem 1.4 with $H_{\textnormal{(ii)}}\leq\frac{1}{2}h_{k}^{*}-1$.
@@
There exists $\xi\in\mathbb{R}$ such that, for every
$\ell\geq h_{0}:=\begin{cases}2^{k-1}&\text{ if }1\leq k\leq 11,\\ \lceil k^{2}(2\log k+\log\log k+2.5)\rceil-2&\text{ if }k\geq 12,\end{cases}$
we have
$\sum_{n\leq x}r_{\mathbb{P}^{k},\ell}(n)^{2}\ll x^{2\ell/k-1}(\log x)^{\xi}.$
@@
$\sum_{n\leq x}r_{\mathbb{P}^{k},\ell}(n)^{2}\leq\int_{0}^{1}|g(\alpha;x)|^{2\ell}\,\mathrm{d}\alpha\leq\sum_{n\leq\ell x}r_{\mathbb{P}^{k},\ell}(n)^{2},$
@@
$n_{1}+\cdots+n_{\ell}=m_{1}+\cdots+m_{\ell}$ (5.8)
with $n_{i},m_{j}\leq x$ and $n_{i},m_{j}\in\mathbb{P}^{k}$.
For $1\leq k\leq 11$, by Hua’s inequality (cf. Hua [8, Theorem 4]), there exists $c(k)\in\mathbb{R}$ such that
@@
$\ll x^{2\ell/k-1}(\log x)^{c(k)}.$
For $k\geq 12$, by [8, Lemma 7.13] we have
$\int_{0}^{1}\bigg{|}\sum_{1\leq n\leq x}e(\alpha n^{k})\bigg{|}^{2\ell}\,\mathrm{d}\alpha\ll x^{2\ell-k}$
for every $\ell>k^{2}(2\log k+\log\log k+2.5)-2$, so the lemma follows by the same argument as above. ∎